if the tangent at (1,-2) to the circle x2+y2=5 touches the circle
x2+y2-8x+6y+20=0 at P.then Py=?

Grade:12

1 Answers

Arun
5 years ago
Tangent at (1, -2) of circle x² + y² = 5
 
x -2y = 5
 
Now
Centre of another circle = (4, -3)
 
Now
x-2y- 5 = 0
 
2x1 + 2y1 -4x -4x1 + 3y  +3y1 +20 = 0
 
 
Now compare and you will find the point.